the visible light of wavelength 350 NM to 800 NM is incident normally on a thin film with refractive index 4 it is observed that three wavelength 720 NM, 480 NM and 360 NM are compulsory missing from reflected spectra find the minimum thickness of a film.
1
Expert's answer
2020-12-08T11:00:19-0500

As per the given question,

Refractive index of the film"(\mu) = 4"

Minimum wavelength "(\lambda) = 360nm"

"2\mu t = (2n-1)\lambda"

Now, substituting the values,

"t= \frac{(2n-1)\lambda}{2\mu}"


"\Rightarrow t= \frac{(2-1)\times 360\times 10^{-9}}{2\times 4}"


"\Rightarrow t = 45\times 10^{-9}m"